OpenStudy (anonymous):

In solving the equation (x + 4)(x – 7) = -24, Eric stated that the solution would be x + 4 = -24 => x = -28 or (x – 7) = -24 => x = -17 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

this sound like it was written by a math teacher. if the product of two numbers is -24 you have no idea what the two numbers are. one is z, the other is \[\frac{-24}{z}\] if you want to solve this quadratic, you must multiply out and collect terms, set = 0 and solve.

OpenStudy (radar):

Yep, Eric messed up, he did not solve the problem.

OpenStudy (anonymous):

\[(x+4)(x-7)=-24\] \[x^2-3x-28=-24\] \[x^2-3x-52=0\] etc

OpenStudy (anonymous):

that is wrong it would be x^2 -3x -4=0

OpenStudy (anonymous):

yes you are right and i am wrong makes it much easier

OpenStudy (anonymous):

lol satellite \[x^2 - 3x - 28 = -24\] \[x^2 - 3x -4 = 0\]